Harry is 1/3 as old as Ron and 8 years younger than Hermoine, find the sum of their ages if Harry is 8 years old
Answer by satyareddy22(84)

let x,y,z be the ages of harry,ron and hermoine respectively.
Harry is 1/3 as old as Ron
=>x=y/3--------------(1)
z-x=8----------------(2)
Harry is 8 years old
so, from (1) 8=y/3 =>y=24 yrs
again from (2)z=8+x=8+8=16=>z=16yrs
The sum of their ages = 8+16+24=48 yrs